Circular No. 7335

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gn IN M61 Alessandro Dimai, Cortina, Italy, reports his discovery of an apparent

supernova (mag 16.0) on an arm of the host galaxy (NGC 4303 = M61), found on a CCD frame taken on Dec. 17.22 UT with a 0.50-m telescope. He reported the new object as being 35′′ east and 25′′ south of the galaxy nucleus; the new star is not present in J. Wray’s Color Atlas of Galaxies. W. Li, University of California at Berkeley, reports that an image taken on Dec. 18.5 with the Katzman Automatic Imaging Telescope (KAIT) shows SN 1999gn at mag ∼ 15.5. Li measured a position for SN 1999gn: α = 12^{h}21^{m}57^{s}.02, δ = +4^{o}27′45′′.6 (equinox 2000.0), which is 31′′.7 east and 39′′.8 south of the nucleus of M61. A KAIT image taken on Nov. 26.5 showed nothing at the position of the new star (limiting mag ∼ 19.0). Previous supernovae in M61 were 1926A, 1961I, and 1964F.

COMET C/1999 S4 (LINEAR) Improved parabolic orbital elements from MPC 37026 (and ephemeris

extension to IAUC 7285):

T = 2000 July 26.1053 TT  ω = 151.^{o}0719
Ω = 83.1520 }2000.0
q = 0.763984 AU  i = 149.3583
1999/2000 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 28  2^{h}21.^{m}84 +33^{o}00.^{′}8 2.627 3.305 126.^{o}2 13.^{o}9 14.3
Jan. 7  2 05.20 +31 22.1 2.673 3.187 112.9 16.5 14.2
17  1 52.35 +29 49.1 2.737 3.067 100.1 18.4 14.1
27  1 43.03 +28 28.4 2.811 2.946  88.0 19.5 13.9
Feb. 6  1 36.81 +27 23.4 2.884 2.823  76.6 19.9 13.8
16  1 33.19 +26 35.5 2.948 2.699  65.8 19.5 13.7
26  1 31.69 +26 03.9 2.997 2.573  55.7 18.5 13.5
Mar. 7  1 31.89 +25 47.6 3.025 2.445  46.1 17.0 13.3
17  1 33.44 +25 45.4 3.028 2.315  37.2 15.1 13.1

Visual m_{1} estimates: Nov. 13.51 UT, 13.9 (Y. Nagai, Yamanashi, Japan, 0.40-m reflector); 17.45, 13.8 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 27.76, 14.0 (W. Hasubick, Buchloe, Germany, 0.44-m reflector); Dec. 7.03, 13.7 (J. Carvajal, Avila, Spain, 0.45-m reflector); 9.88, 14.0 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

1999 December 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green